### v5

@@ Abstract
Clustering of nodes in Bayesian Networks (BNs) and related graphical models such as Dynamic BNs (DBNs) has been demonstrated to enhance computational efficiency and improve model learning. Typically, it involves the partitioning of the underlying Directed Acyclic Graph (DAG) into cliques, or optimising for some cost or criteria. Computational cost is important since BN and DBN inference, such as estimating marginal distributions given evidence or updating model parameters, is NP-hard. The challenge is exacerbated by cost dependency, where inference outcomes and hence clustering cost depends on both nodes within a cluster and the mapping of clusters that are connected by at least one arc.
We propose an algorithm called Dependent Cluster MAPping (DCMAP) which can, given an arbitrarily defined, positive cost function, iteratively and rapidly find near-optimal, then optimal cluster mappings. DCMAP is shown analytically to be optimal in terms of finding all of the least cost cluster mapping solutions and with no more iterations than an equally informed algorithm. DCMAP is demonstrated on a complex systems seagrass DBN, which has 25 nodes per time-slice, and captures biological, ecological and environmental dynamics and their interactions to predict the impact of dredging stressors on resilience and their cumulative effects over time. The algorithm is employed to find clusters to optimise the computational efficiency of inferring marginal distributions given evidence. For the 25 (one time-slice) and 50-node (two time-slices) DBN, the search space size was $9.91\times 10^{9}$ and $1.51\times 10^{21}$ possible cluster mappings, respectively, but the first optimal solution was found at iteration number 856 (95% CI 852,866), and 1569 (1566,1581) with a cost that was 4% and 0.2% of the naive heuristic cost, respectively. Through optimal clustering, DCMAP opens up opportunities for further research beyond improving computational efficiency, such as using clustering to minimise entropy in BN learning.
keywords:
bayesian network, clustering, optimisation, dynamic programming, entropy
1 Introduction
@@
Bayesian Networks (BNs) and related methods including Dynamics BNs (DBNs) are a form of graphical model that are represented as Directed Acyclic Graphs (DAGs), where relationships between nodes (factors) are shown as arcs and quantified with conditional probabilities [22]. They have been applied widely across domains including medicine, environment, economics and social sciences to provide explanatory decision support, integrating data and even expert knowledge [27, 11]. However, one of the key challenges associated with inference and learning of BNs and DBNs is computation, which is NP-hard [21, 19]. To this end, different approaches to clustering of BN nodes have been used in the context of supporting more computationally efficient inference of marginal distributions given evidence, and model learning, to update the structure of the graph and/or the conditional probabilities associated with each node [5, 15, 12, 18].
@@
Optimal clustering of a BN seeks to minimise some cost or objective function associated with the mapping of BN nodes to clusters. Typically, each node is uniquely mapped to a cluster and the total cost is the sum of costs associated with each cluster. However, a key challenge is cost dependency. Consider a BN $H(\boldsymbol{X},\boldsymbol{A})$ with nodes $X_{i}\in\boldsymbol{X}$ and arcs (or edges) $\boldsymbol{A}$ that encode conditional dependence between parent and child nodes. Fundamentally, BN inference is underpinned by Bayes theorem where the joint probability distribution over all nodes $\boldsymbol{X}$ and given observations (i.e. evidence $\boldsymbol{E}$) is [15]:
$P(\boldsymbol{X},\boldsymbol{E})$ $=\prod_{\forall X_{j}\in\boldsymbol{X}}P(X_{j}|Par(X_{j}))\delta(X_{j})$ (1)
@@
where $P(X_{j}|Par(X_{j}))$ is the conditional probability of node $X_{j}$ given its parents, and $\delta(X_{j})$ the evidence. For a discrete node with virtual evidence, $\delta(X_{j})=\{P(e|x_{j1}),\ldots,P(e|x_{jm})\}$ for binary event $e$ (observe this evidence or not) and $m$ node states $x_{j1},\ldots,x_{jm}$ [22]. The joint probability distribution arises from the multiplication of conditional probability distributions, each of which is potentially multi-dimensional, comprising a node (or variable) $X_{j}$ and parent nodes $Par(X_{j})$. Hence, intuitively: (i) clustering determines which conditional probability distributions are computed together, impacting computational complexity, and (ii) the computed outcomes of a given cluster are dependent on clusters connected by at least one arc via multiplication as per equation (1). This propagation of partial inference outcomes between clusters has been employed in approaches to BN and DBN inference such as Lauritzen and Spiegelhalter [15] and Wu et al. [26]. The dependency of the local cluster cost on the mapping of connected clusters is referred to as cost dependency, noting that the objective is to minimise the global cost which is a combination (typically additive) of local costs. However, finding optimal clusters (i.e. partitions) is NP-hard in general [4] and cost dependency exacerbates this complexity.
@@
The clustering problem can be formalised as one of assigning a unique integer cluster label $u(X_{j})=k,k=1,\ldots,m$ to each node $X_{j}\in\boldsymbol{X}$ of BN $H(\boldsymbol{X},\boldsymbol{A})$; the total number of clusters $m$ is determined through optimisation. Let $G\bigl{(}H(\boldsymbol{X},\boldsymbol{A}),u(\boldsymbol{X})\bigr{)}$ be the total cost to be minimised through cluster mapping $u(\boldsymbol{X})$ and $G^{\ast}\bigl{(}H(\boldsymbol{X},\boldsymbol{A})\bigr{)}$ the optimal cost, and assume that the total cost is the sum of individual cluster costs $G_{k}$:
$G^{\ast}\bigl{(}H(\boldsymbol{X},\boldsymbol{A})\bigr{)}=\operatorname*{arg\,min}_{u(\boldsymbol{X})}\Bigl{(}G\bigl{(}H(\boldsymbol{X},\boldsymbol{A}),u(\boldsymbol{X})\bigr{)}\Bigr{)}=\operatorname*{arg\,min}_{u(\boldsymbol{X})}\left(\sum_{k=1}^{m}{G_{k}}\right).$ (2)
@@ 1.1 Case Study
We use as a case study a DBN of a seagrass ecosystem [26]. Seagrasses are a critical primary habitat for fish and many endangered species including the dugong and green turtle, and contribute USD\$1.9 trillion per annum through nutrient uptake and cycling. However, better management of human stressors such as dredging requires the modelling of cumulative effects arising from interactions between biological, ecological and environmental factors. The non-homogeneous DBN itself is complex and computationally intensive, and inference is infeasible within a standard (MCMC) framework for real-world decision support contexts. Furthermore, understanding areas of uncertainty in the system and designing data collection strategies to mitigate uncertainty is of utmost interest in managing such complex systems. This uncertainty can be characterised by entropy with clusters corresponding to scenarios of evidence (i.e. observed data) (Section 5.2). We focus on computation as a cost function $G$ to demonstrate how the proposed algorithm can find clusters to minimise the computational complexity of inference, and discuss how our algorithm could also be used to minimise other cost functions such as entropy.
1.2 Existing Work
Clustering of BNs has been employed in a wide range of contexts, evolving from a focus on computation of marginal distributions given evidence such as the seminal work of Lauritzen and Spiegelhalter [15] to predominantly causal discovery and graph learning [11, 18] in recent times. Although the approach of Lauritzen and Spiegelhalter [15] and related approaches such as Kschischang et al. [14] (see Daly et al. [5] for a review) employ clustering, they are based on a single mapping of nodes to clusters (referred to as cliques) using heuristics. They do not evaluate different cluster mappings on an outcome score. These challenges are addressed in more recent literature through three main approaches: (i) constraint based, that evaluate conditional independence structures i.e. if an arc exists or not, then its direction, (ii) scoring, often with heuristics to evaluate goodness of fit of a proposed network, and (iii) hybrid, combining constraints with heuristics [10] .
@@ 1.2 Existing Work
A number of approaches use clusters as a structure to restrict or constrain the local search space in a graph learning framework such as the seminal work of Friedman et al. [7]. They define clusters in a similar vein to clique-trees [15] in a “restrict and maximise” framework to optimise a mutual-information based score of the learned network. Kojima et al. [12] advance this approach, effectively relaxing the cluster tree requirement in Friedman et al. [7] to a cluster graph to help address the issue of large clusters, which make local optimisation computationally infeasible. They used ancestral constraints to help limit the search space of clusters and the algorithm was shown to outperform greedy algorithms including Max-Min Hill Climbing (MMHC) and Hill Climbing (HC) with Tabu search. However, computational feasibility becomes a problem again when there are too many edges between clusters [12].
@@
Lu et al. [18] also adapt the work of Friedman et al. [7] for undirected graphs, using clusters based on proposed parents and children of a node $X$ to constrain the search in an iterative framework based on A*, a heuristic-informed search algorithm. They show empirically that for small networks, exhaustive and in some cases optimal methods such as A* and SAT (Boolean SATisfiability) outperform greedy methods such as PC (Peter and Clark), which can suffer from convergence to local optimal, and their adaptation of A* performs as well as greedy methods in larger networks. However, their approach is exponential in complexity with neighbourhood size (i.e. number of parents).
@@
Zhang et al. [28] present an alternate approach in which clusters are determined through agglomerative clustering using Pearson correlation as the distance metric; however, little is provided in terms of algorithm performance and its convergence characteristics. Huang and Zhou [10] and Gu and Zhou [9] both develop comprehensive approaches that use similar agglomerative clustering based on mutual information and correlation as the distance metric, respectively. The former learns skeletons (an undirected version of the graph) for each cluster using their variant of the PC algorithm and log-likelihood ratio tests of conditional independence. Gu and Zhou [9] use a score-based algorithm to learn partial DAGs for each cluster, which are then fused in the final step proposing edges between clusters. Huang and Zhou [10] combine clusters with automated parameter tuning for turning skeletons into partial then fully directed acyclic graphs and hybrid greedy initialisation, obtaining empirical improvements in computational efficiency and more accurate DAGs. However, in both cases, the performance is dependent on the quality and size of the clusters.
@@
Clearly, clustering has demonstrated utility to improve computational feasibility in learning BNs of varying degrees of optimality given a scoring function (i.e. cost function). However, all of the reviewed methods above focused on scores that changed given changes to the BN vis a vis inclusion or exclusion of edges, rather than scores that changed given different mappings of nodes to clusters. This paper focuses on the latter task of explicit optimisation with respect to clusters, to develop a more generalisable algorithm that works with arbitrary non-negative objective functions. Nevertheless, the literature does highlight the utility of dynamic programming and clustering for optimisation [7, 12, 18].
@@
Note that there also exists substantial literature in graph partitioning such as repeated graph bisection, Kernighan and Lin algorithm, spectral bisection and multilevel partitioning [4, 1]. Zheng et al. [29], for instance, demonstrate how dynamic programming can be used to optimally and efficiently cluster nodes in the absence of cost dependency, applied to convolutional neural networks. However, the problem context for these approaches typically involves partitioning a graph into a set number of clusters or to satisfy some criteria for each cluster defined with respect to static node and/or edge weights. This differs significantly from the BN context and associated cost dependency between clusters.
1.3 Layout of Paper
@@
We develop a novel algorithm called DCMAP (Dependent Cluster MAPping) that optimises BN cluster mappings given an arbitrary, non-negative scoring function which we refer to as a cost function in the presence of cost dependency. Such dependency arises in the BN context of using clustering to improve computational efficiency when inferring marginal distributions given evidence [15] and in model learning [7, 18]. Noting as background the context of BN inference (Section 2.1) and introducing an illustrative example (Section 2.2), we first present an intuitive overview of DCMAP and its notation (Section 2.3), followed by its derivation from dynamic programming principles (Section 2.3.1 and 2.3.3). The algorithm is presented in detail in Section 3 and shown analytically to find all the optimal solutions such that an equally informed algorithm cannot do so in fewer iterations. An empirical demonstration is provided on the case study usage scenario of using clusters to minimise computation cost of inference for our complex systems seagrass DBN in Section 4. Reflections and suggestions for future work, including extension to other cost functions for optimisation such as entropy are provided in the Discussion (Section 5).
2 DCMAP Algorithm: Background and Overview
@@ 2.1 Background: BN Inference
$P(X,\boldsymbol{E})$ $=\sum_{\forall X_{i}\in\boldsymbol{X}\backslash X}\prod_{\forall X_{i}\in\boldsymbol{X}}P(X_{i}|Par(X_{i}))\delta(X_{i})$ (3)
[…]
2.2 Illustrative Example

### v6

@@ Abstract.
Background: Clustering of nodes in Bayesian Networks (BNs) and related graphical models such as Dynamic BNs (DBNs) has been demonstrated to enhance computational efficiency and improve model learning. It typically involves partitioning the underlying Directed Acyclic Graph (DAG) into cliques or optimising for some cost or criteria.
Objectives: We focus on a critical but understudied aspect of optimal clustering involving cost dependency. This is where inference outcomes and hence clustering costs depend on both nodes within a cluster and the mapping of clusters that are connected by at least one arc.
Methods: We propose a novel algorithm called Dependent Cluster MAPping (DCMAP) which can, given an arbitrary, positive cost function, iteratively and rapidly find near-optimal, then optimal cluster mappings.
Results: DCMAP is shown analytically to be optimal in terms of finding all of the least cost cluster mapping solutions and with no more iterations than an equally informed algorithm. Demonstrated on a complex systems seagrass DBN with $9.91\times 10^{9}$ and $1.51\times 10^{21}$ possible cluster mappings for 25 and 50 node configurations, it took 856 and 1569 iterations on average to find the first optimal solution, respectively.
Conclusions: The effectiveness of DCMAP enables future research in BN learning using optimisation, such as through enhancing computational efficiency or minimising entropy for learning. This is critically important as computation of marginal distributions or updating model parameters is NP-hard.
copyright: ccdoi: 10.1613/jair.1.xxxxxjournalvolume: 4article: 6publicationmonth: 8journalyear: 2025
1. Introduction
@@
Bayesian Networks (BNs) and related methods including Dynamics BNs (DBNs) are a form of graphical model that are represented as Directed Acyclic Graphs (DAGs), where relationships between nodes (factors) are shown as arcs and quantified with conditional probabilities (Pearl1988). They have been applied widely across domains including medicine, environment, economics and social sciences to provide explanatory decision support, integrating data and even expert knowledge (Wu2017; Kitson2023). However, one of the key challenges associated with inference and learning of BNs and DBNs is computation, which is NP-hard (Park2004; Murphy2002). To this end, different approaches to clustering of BN nodes have been used in the context of supporting more computationally efficient inference of marginal distributions given evidence, and model learning, to update the structure of the graph and/or the conditional probabilities associated with each node (Daly2011; Lauritzen1988; Kojima2010; Lu2021). Appendix A illustrates with a simple example how three different clustering approaches reduce computational cost to different degrees. A general approach to optimisation of BN clusters (i.e. with aribtrary criteria) can advance research in probabilistic graphical models by providing a benchmark for validating approximate inference methods, which are necessary in high-dimensional problems, and enable new methods for computation and inference such as by finding homogeneous or correlated regions in a non-homogeneous system (Lin2020; Wu2018).
@@
Optimal clustering of a BN seeks to minimise some cost or objective function associated with the mapping of BN nodes to clusters. Typically, each node is uniquely mapped to a cluster and the total cost is the sum of costs associated with each cluster. However, a key challenge is cost dependency. Consider a BN $H(\boldsymbol{X},\boldsymbol{A})$ with nodes $X_{i}\in\boldsymbol{X}$ and arcs (or edges) $\boldsymbol{A}$ that encode conditional dependence between parent and child nodes. Fundamentally, BN inference is underpinned by Bayes theorem where the joint probability distribution over all nodes $\boldsymbol{X}$ and given observations (i.e. evidence $\boldsymbol{E}$) is (Lauritzen1988):
(1) $P(\boldsymbol{X},\boldsymbol{E})$ $=\prod_{\forall X_{j}\in\boldsymbol{X}}P(X_{j}|Par(X_{j}))\delta(X_{j})$
@@
where $P(X_{j}|Par(X_{j}))$ is the conditional probability of node $X_{j}$ given its parents, and $\delta(X_{j})$ the evidence, which can be the inputs to or observations for inference. The joint probability distribution arises from the multiplication of conditional probability distributions, each of which is potentially multi-dimensional, comprising a node (or variable) $X_{j}$ and parent nodes $Par(X_{j})$. Hence, intuitively: (i) clustering determines which conditional probability distributions are computed together, impacting computational complexity, and (ii) the computed outcomes of a given cluster are dependent on clusters connected by at least one arc via multiplication as per equation (1). This propagation of partial inference outcomes between clusters has been employed in approaches to BN and DBN inference such as (Lauritzen1988) and (Wu2018). The dependency of the local cluster cost on the mapping of connected clusters is referred to as cost dependency, noting that the objective is to minimise the global cost which is a combination (typically additive) of local costs. However, finding optimal clusters (i.e. partitions) is NP-hard in general (Buluc2016) and cost dependency exacerbates this complexity.
@@
The clustering problem can be formalised as one of assigning a unique integer cluster label $u(X_{j})=k,k=1,\ldots,m$ to each node $X_{j}\in\boldsymbol{X}$ of a BN with DAG $H(\boldsymbol{X},\boldsymbol{A})$ and conditional probability parameters $\boldsymbol{\theta}=\{P(X_{j}|Par(X_{j})\}$; the total number of clusters $m$ is determined through optimisation. Since it is the DAG nodes that are clustered, for notational simplicity, let $G\bigl(H(\boldsymbol{X},\boldsymbol{A}),u(\boldsymbol{X})\bigr)$ be the total cost to be minimised through cluster mapping $u(\boldsymbol{X})$, implicitly capturing $\boldsymbol{\theta}$ through the DAG. Additionally, $G^{\ast}\bigl(H(\boldsymbol{X},\boldsymbol{A})\bigr)$ is the optimal cost, and assume that the total cost is the sum of individual cluster costs $G_{k}$:
(2) $G^{\ast}\bigl(H(\boldsymbol{X},\boldsymbol{A})\bigr)=\operatorname*{arg\,min}_{u(\boldsymbol{X})}\Bigl(G\bigl(H(\boldsymbol{X},\boldsymbol{A}),u(\boldsymbol{X})\bigr)\Bigr)=\operatorname*{arg\,min}_{u(\boldsymbol{X})}\left(\sum_{k=1}^{m}{G_{k}}\right).$
@@ 1.1. Case Study
We use as a case study a DBN of a seagrass ecosystem (Wu2018) to show how clustering can minimise computational cost. Seagrasses are a critical primary habitat for fish and many endangered species including the dugong and green turtle, and contribute USD\$1.9 trillion per annum through nutrient uptake and cycling. However, better management of human stressors such as dredging requires the modelling of cumulative effects arising from interactions between biological, ecological and environmental factors. The non-homogeneous DBN itself is complex and computationally intensive, and inference is infeasible within a standard (MCMC) framework for real-world decision support contexts. Furthermore, understanding areas of uncertainty in the system and designing data collection strategies to mitigate uncertainty is of utmost interest in managing such complex systems. This uncertainty can be characterised by entropy with clusters corresponding to scenarios of evidence (i.e. observed data) (Section 5.2). We focus on computation as a cost function $G$ to demonstrate how the proposed algorithm can find clusters to minimise the computational complexity of inference, and discuss how our algorithm could also be used to minimise other cost functions such as entropy.
1.2. Existing Work
Clustering of BNs has been employed in a wide range of contexts, evolving from a focus on computation of marginal distributions given evidence such as the seminal work of (Lauritzen1988) to predominantly causal discovery and graph learning (Kitson2023; Lu2021) in recent times. Almost all existing methods for inferring marginal distributions use heuristics based on the graph to form clusters, such as triangulation in join-tree inference (Lauritzen1988), conditioning sets (cutsets) in bucket elimination (Dechter1999), factor graphs (Kschischang2001) and approximate inference with cutsets (Bidyuk2007). They help tackle the challenge of exponential growth in computational complexity with induced tree width (Lin2020). In some methods, such as mini-clustering (Mateescu2010), incremental region selection (Forouzan2015), clusters for selective filtering in a DBN (Albrecht2016), and triplet region construction (Lin2020), the heuristically formed clusters are refined through greedy search or local optimisation based on some function such as conditional entropy. Much of the more recent work focuses on approximate methods to tackle high-dimensional problems. However, explicit study of global optimisation of clusters and their impact on computation and inference is missing in this literature.
In contrast, the BN learning literature has seen some progress in clustering and inference towards global optimisation. There are three main approaches: (i) constraint based, that evaluate conditional independence structures i.e. if an arc exists or not, then its direction, (ii) scoring, often with heuristics to evaluate goodness of fit of a proposed network, and (iii) hybrid, combining constraints with heuristics (Huang2022) .
@@ 1.2. Existing Work
A number of approaches use clusters as a structure to restrict or constrain the local search space in a graph learning framework such as the seminal work of (Friedman1999). They define clusters in a similar vein to clique-trees (Lauritzen1988) in a “restrict and maximise” framework to optimise a mutual-information based score of the learned network. (Kojima2010) advance this approach, effectively relaxing the cluster tree requirement in (Friedman1999) to a cluster graph to help address the issue of large clusters, which make local optimisation computationally infeasible. They used ancestral constraints to help limit the search space of clusters and the algorithm was shown to outperform greedy algorithms including Max-Min Hill Climbing (MMHC) and Hill Climbing (HC) with Tabu search. However, computational feasibility becomes a problem again when there are too many edges between clusters (Kojima2010).
@@
(Lu2021) also adapt the work of (Friedman1999) for undirected graphs, using clusters based on proposed parents and children of a node $X$ to constrain the search in an iterative framework based on A*, a heuristic-informed search algorithm. They show empirically that for small networks, exhaustive and in some cases optimal methods such as A* and SAT (Boolean SATisfiability) outperform greedy methods such as PC (Peter and Clark), which can suffer from convergence to local optimal, and their adaptation of A* performs as well as greedy methods in larger networks. However, their approach is exponential in complexity with neighbourhood size (i.e. number of parents).
@@
(Zhang2018) present an alternate approach in which clusters are determined through agglomerative clustering using Pearson correlation as the distance metric; however, little is provided in terms of algorithm performance and its convergence characteristics. (Huang2022) and (Gu2020) both develop comprehensive approaches that use similar agglomerative clustering based on mutual information and correlation as the distance metric, respectively. The former learns skeletons (an undirected version of the graph) for each cluster using their variant of the PC algorithm and log-likelihood ratio tests of conditional independence. (Gu2020) use a score-based algorithm to learn partial DAGs for each cluster, which are then fused in the final step proposing edges between clusters. (Huang2022) combine clusters with automated parameter tuning for turning skeletons into partial then fully directed acyclic graphs and hybrid greedy initialisation, obtaining empirical improvements in computational efficiency and more accurate DAGs. However, in both cases, the performance is dependent on the quality and size of the clusters.
@@
Clearly, clustering has demonstrated utility to improve computational feasibility in BN inference of marginal distributions and learning BNs with varying degrees of optimality given a scoring function (i.e. cost function). However, none of the reviewed literature explicitly studied globally optimal clustering and its impact on inference and computation. Existing BN learning methods focus on inclusion or exclusion of edges, rather than scores that change given different mappings of nodes to clusters. Given the pervasive role clustering plays in BN research and the utility of dynamic programming and clustering for optimisation (Friedman1999; Kojima2010; Lu2021), this paper focuses on explicit, generalisable optimisation of clusters in the context of BN inference.
@@
Note that there also exists substantial literature in graph partitioning such as repeated graph bisection, Kernighan and Lin algorithm, spectral bisection and multilevel partitioning (Buluc2016; Bader2013). (Zheng2020), for instance, demonstrate how dynamic programming can be used to optimally and efficiently cluster nodes in the absence of cost dependency, applied to convolutional neural networks. However, the problem context for these approaches typically involves partitioning a graph into a set number of clusters or to satisfy some criteria for each cluster defined with respect to static node and/or edge weights. This differs significantly from the BN context and associated cost dependency between clusters.
1.3. Layout of Paper
@@
We develop a novel algorithm called DCMAP (Dependent Cluster MAPping) that optimises BN cluster mappings given an arbitrary, non-negative scoring function which we refer to as a cost function in the presence of cost dependency. Such dependency arises in the BN context of using clustering to improve computational efficiency when inferring marginal distributions given evidence (Lauritzen1988) and in model learning (Friedman1999; Lu2021). A simple example is provided in Appendix A showing how DCMAP (least computation) and two other popular clustering approaches reduce computation. Noting as background the context of BN inference (Section 2.1) and introducing an illustrative example (Section 2.2), we first present an intuitive overview of DCMAP and its notation (Section 2.3), followed by its derivation from dynamic programming principles (Section 2.3.1 and 2.3.3). The algorithm is presented in detail in Section 3 and shown analytically to find all the optimal solutions such that an equally informed algorithm cannot do so in fewer iterations. An empirical demonstration is provided on the case study usage scenario of using clusters to minimise computation cost of inference for our complex systems seagrass DBN in Section 4. Reflections and suggestions for future work, including extension to other cost functions for optimisation such as entropy are provided in the Discussion (Section 5).
2. DCMAP Algorithm: Background and Overview
@@ 2.1. Background: BN Inference
(3) $P(X,\boldsymbol{E})$ $=\sum_{\forall X_{i}\in\boldsymbol{X}\backslash X}\prod_{\forall X_{i}\in\boldsymbol{X}}P(X_{i}|Par(X_{i}))\delta(X_{i})$
where $P(X_{i}|Par(X_{i}))$ is the conditional probability of node $X_{i}$ given its parents, $\backslash$ denotes set difference, and $\delta(X_{i})$ is the evidence for node $X_{i}$. Evidence can be used to, among other things, incorporate observations and evaluate scenarios. The joint distribution (1) has $|\boldsymbol{X}|$ nodes or dimensions, and $\sum_{X_{i}}$ denotes a summation over all discrete states of node $X_{i}$ which we denote as lowercase $x_{i1},x_{i2},\ldots$, effectively removing that dimension from the joint distribution through marginalisation. Note that $P(\boldsymbol{E})$ is a constant that normalises the resulting distribution (Lauritzen1988). Although the notation focuses on discrete probability distributions to minimise complexity, the proposed clustering algorithm is generalisable to continuous distributions since it ultimately relies only on an arbitrary non-negative cost function.
2.2. Illustrative Example